OpenStudy (anonymous):

What is the exact value of the expression 11sqrt(8)+6sqrt(12)−5sqrt(2)? Show your work.

OpenStudy (cwrw238):

11sqrt8 + 6sqrt12 - 5sqrt2 =11 sqrt4sqrt2 + 6sqrt4sqrt3 - 5sqrt2 = 22sqrt2 + 12sqrt3 - 5sqrt2 = 17sqrt2 + 12 sqrt3

OpenStudy (goformit100):